### v2

# Efficient Chromatic-Number-Based Multi-Qubit Decoherence and Crosstalk Suppression

Abstract
The performance of quantum computers is hindered by decoherence and crosstalk, which cause errors and limit the ability to perform long computations. Dynamical decoupling is a technique that alleviates these issues by applying carefully timed pulses to individual qubits, effectively suppressing unwanted interactions. However, as quantum devices grow in size, it becomes increasingly important to minimize the time required to implement dynamical decoupling across the entire system. Here, we present “Chromatic-Hadamard Dynamical Decoupling” (CHaDD), an approach that efficiently schedules dynamical decoupling pulses for quantum devices with arbitrary qubit connectivity. By leveraging Hadamard matrices, CHaDD achieves a circuit depth that scales linearly with the chromatic number of the connectivity graph for general two-qubit interactions, assuming instantaneous pulses. This includes $ZZ$ crosstalk, which is prevalent in superconducting qubit devices. CHaDD’s scaling represents an exponential improvement over all previous multi-qubit decoupling schemes for devices with connectivity graphs whose chromatic number grows at most polylogarithmically with the number of qubits. For graphs with a constant chromatic number, CHaDD’s scaling is independent of the number of qubits. Our results suggest that CHaDD can become a useful tool for enhancing the performance and scalability of quantum computers by efficiently suppressing decoherence and crosstalk across large qubit arrays.
Introduction.—Quantum computers contain interconnected qubits that experience decoherence and control errors due to spurious interactions with the environment and surrounding qubits. This adversely affects such devices’ ability to retain or process information for periods of time longer than the decoherence timescale, which is necessary for achieving a quantum advantage over classical computers [1]. In order to surmount this obstacle, considerable attention has been devoted to developing methods that mitigate these deleterious effects, among which dynamical decoupling (DD) has recently been shown to play an important role. DD is an error mitigation technique that averages out undesired Hamiltonian terms by applying deterministic [2, 3, 4, 5] or random sequences of scheduled pulses [6, 7]. Originally designed to suppress low-frequency noise in nuclear magnetic resonance [8, 9, 10, 11], DD effectively decouples the quantum system from both external noise due to decoherence [12, 13, 14, 15, 16] and internal noise due to undesired always-on interactions [17, 18, 19, 20, 21, 22]. DD has recently found numerous applications in quantum information processing, e.g., in noise characterization [23, 24, 25, 26, 27, 28] and improving algorithmic performance [29, 30, 31, 32, 33, 34].
Most of the attention in improving performance through DD has focused on the development of DD sequences that achieve high perturbation-theory-order noise cancellation [35, 36, 37, 38, 39, 40, 41, 42, 43, 44], robustness to pulse errors [45, 46, 47, 48], or reduction of the requirements for quantum error correction [49, 50, 51, 52, 53] and error avoidance [54, 55, 56]. Much less attention has been paid to developing efficient DD pulse sequences for simultaneously decoupling multiple interconnected qubits, and studies of this topic date primarily to early results using Hadamard matrices and orthogonal arrays [17, 57, 58, 59, 60, 61, 62], along with more recent developments [63, 18, 20, 21, 22]. This problem grows in complexity for qubit layouts that are more highly connected, i.e., as the qubit graph degree grows. It is generally recognized that a higher graph degree is desirable [64, 65], as this reduces the overhead associated with coupling geometrically distant qubits.
In this work, we introduce Chromatic-Hadamard Dynamical Decoupling (CHaDD), which provides an efficient solution to completely decoupling an arbitrary qubit interaction graph $G=(V,E)$, where $V$ is the vertex (i.e., qubit) set and $E$ the edge (i.e., qubit-qubit coupling) set. We measure efficiency in terms of circuit depth, i.e., the number of time steps. This is equivalent to the number of applied pulses, where simultaneous pulses are counted as a single pulse. Until recently, it was thought that efficient schemes should be at most linear in the total number of qubits $n=|V|$ [60, 61, 62] for instantaneous (“bang-bang” [2]) pulses. Then, it was established that scaling of complete decoupling of a $ZZ$ crosstalk graph [66] could be stated in terms of the chromatic number of the graph $\chi(G)$ (an idea dating back to Refs. [17, 18]), i.e., the minimum number of distinct colors required to properly color the graph, instead of the number of qubits, albeit with exponential scaling, $2^{\chi(G)}$ [22].
Here, we show that it is possible to achieve efficient, universal, first-order “bang-bang” decoupling. Namely, we prove that in general, the circuit depth scales linearly with the chromatic number, with an improved prefactor for single-axis decoupling (e.g., the special case of $ZZ$ crosstalk). This is an exponential improvement over Ref. [22]’s single-axis crosstalk suppression result, as well as over all previous multi-axis decoupling schemes building on Hadamard matrices or orthogonal arrays [57, 58, 59, 60, 61, 62] for graphs with a chromatic number that scales at most polylogarithmically in $n$. The schemes of Refs. [60, 61, 62] achieve parity with CHaDD when $\chi(G)\sim n$, such as trapped-ion devices with all-to-all connectivity [64]. In devices with a constant chromatic number, such as all current superconducting quantum computing hardware, CHaDD’s scaling is independent of $n$, providing a significant advantage over decoupling schemes that scale linearly with the number of qubits. In such cases, CHaDD can further help diagnose and suppress crosstalk between non-natively coupled qubits corresponding to an effective connectivity graph with a higher chromatic number.
Graph coloring.—Coloring is the assignment of labels (often called colors) to the vertices of a graph such that no two adjacent vertices share the same color. The chromatic number is related to the graph degree: Brooks’ Theorem states that for a connected graph $G$ with maximum degree $\Delta$, $\chi(G)\leq\Delta$ except when $G$ is a complete graph or an odd cycle, in which case $\chi(G)=\Delta+1$ [67, 68]. Moreover, if the maximum degree of the subgraph induced on a vertex $v\in V$ and its neighborhood is $\Delta(v)$, then $\chi(G)\geq\Delta(v)+1$, which provides a lower bound using neighborhood degree [69]. Lower bounds also exist in terms of the adjacency matrix [70, 71]. Finding the chromatic number is one of Karp’s $21$ NP-complete problems [72], but in many cases of interest, we do not need to exactly know $\chi(G)$: since $\Delta$ is often a constant in the quantum computing context, we may in such cases simply replace $\chi(G)$ by the graph degree $\Delta$ to obtain an upper bound.
Error model.—Consider a system of qubits occupying the vertices $V$ of a graph $G$. We are concerned with the suppression of undesired interactions between this system and its environment and the suppression of undesired internal system terms arising, e.g., due to crosstalk. Letting $\{\sigma_{v}^{\alpha}\}_{\alpha\in\{x,y,z\}}$ denote the Pauli matrices acting only on qubit $v$, quite generally this scenario can be represented in terms of the following Hamiltonian:
$H$ $=H_{1}+H_{2},$ (1a)
$H_{1}$ $=\sum_{\alpha\in\{x,y,z\}}H_{1}^{\alpha}\ ,\ \ H_{2}=\sum_{\alpha,\beta\in\{x,y,z\}}H_{2}^{\alpha\beta},$ (1b)
$H_{1}^{\alpha}$ $=\sum_{v\in V}\sigma^{\alpha}_{v}\otimes B^{\alpha}_{v}\ ,\ \ H_{2}^{\alpha\beta}=\!\!\!\sum_{\{(u,v)\in E\ |\ u<v\}}\!\!\!\sigma^{\alpha}_{u}\sigma^{\beta}_{v}\otimes B^{\alpha\beta}_{uv}.$ (1c)
In Eq. 1c, we sum over ordered pairs of vertices, and $B^{\alpha}_{v}=\omega^{\alpha}_{v}\tilde{B}^{\alpha}_{v}$ ($B^{\alpha\beta}_{uv}=J^{\alpha\beta}_{uv}\tilde{B}^{\alpha\beta}_{uv}$), where $\tilde{B}^{\alpha}_{v}$ ($\tilde{B}^{\alpha\beta}_{uv}$) are dimensionless operators that act purely on the environment, and $\omega^{\alpha}_{v}$ ($J^{\alpha\beta}_{uv})$ are the corresponding couplings with dimensions of energy. In the case of undesired internal terms, the $\tilde{B}^{\alpha}_{v}$ ($\tilde{B}^{\alpha\beta}_{uv}$) represent the identity operator on all qubits but $v$ (and $u$). In the former case, the unperturbed evolution given by the “free evolution unitary” $f_{\tau}=\exp(-i\tau H)$ is generally non-unitary and decoherent, while in the latter it is unitary but subject to coherent errors.
Dynamical Decoupling.—DD is based on the application of short and narrow (“bang-bang” [2]) pulses $P_{\alpha,v}(\theta)=\exp[-i(\theta/2)\sigma_{v}^{\alpha}]$. Here we are concerned primarily with the case of $\pi$-pulses and denote $X_{v}=P_{x,v}(\pi)=-i\sigma_{v}^{x}$ (an “$X$ pulse”), and similarly for $Y$ and $Z$. It is well known that in order to suppress decoherence and coherent errors, it suffices to apply $\pi$ pulses at regular intervals, as long as these pulses cycle over the elements of a group [4]. If one were to try to apply this same sequence synchronously to all qubits in an attempt to suppress both $H_{1}$ and $H_{2}$, one would accomplish the former but inadvertently restore the latter. To dynamically decouple several qubits in a manner that suppresses both $H_{1}$ and $H_{2}$, one can use Hadamard matrices and orthogonal arrays to schedule the pulses in such a way that they do not restore the undesired interactions  [57, 58, 59, 60, 61, 62]. However, this does not account for the qubit connectivity graph, and recent work that does resulted in a scheme requiring a cost of $2^{\chi(G)}$ pulses to decouple all $ZZ$ crosstalk in a graph $G$ [22] (the special case $H=H_{2}^{zz}$). We now show how this can be both exponentially reduced in cost and extended to deal with a general Hamiltonian $H$ [Eq. 1].
Results.—We first specialize to the case where $H_{1}$ and $H_{2}$ do not include pure-$x$-type terms. This includes $ZZ$ crosstalk.
Theorem 1 (single-axis CHaDD).
Assume the free evolution unitary is $f_{\tau}=e^{-i\tau H}$, where $H$ is given by Eq. 1. Then a circuit depth of $\chi\leq N\leq 2\chi$, involving only $X$ pulses, suffices to cancel to first order in $\tau$ all terms in $H$ excluding $H_{1}^{x}$ and $H_{2}^{xx}$ on a qubit connectivity graph $G$ with chromatic number $\chi$.
Proof.
Let $\chi$ be the chromatic number of the graph $G$ corresponding to the Hamiltonian $H$ given by Eq. 1, and let $f:V\rightarrow\left\{1,2,\dots,\chi(G)\right\}\equiv C$ be a proper $\chi$-coloring of the graph $G$. Define $V_{c}\equiv\{v\in V\ |\ f(v)=c\}$, i.e, the set of all vertices of the same color $c$, and $E_{c_{1},c_{2}}\equiv\{(u,v)\in E\ |\ u<v,f(u)=c_{1},f(v)=c_{2}\}$, i.e., the set of all ordered pairs of vertices of colors $c_{1}$ and $c_{2}$, respectively, joined by edges in $E$. Now we return to Eq. 1c and group the terms according to $G$’s coloring:
$H^{\alpha}_{1}$ $=\sum_{c\in C}\sum_{v\in V_{c}}\sigma^{\alpha}_{v}\ ,\ \ \ H^{\alpha\beta}_{2}=\sum_{c_{1}\neq c_{2}}\sum_{(u,v)\in E_{c_{1},c_{2}}}\sigma^{\alpha}_{u}\sigma^{\beta}_{v}.$ (2)
We have suppressed the bath operators in Eq. 2 for notational simplicity; they will not matter in our calculations below since we will only consider first-order time-dependent perturbation theory through the Baker-Campbell-Hausdorff (BCH) expansion; the effect of non-commuting bath operators appears only to second order.
If we conjugate the free evolution unitary $f_{\tau}$ by $X$ pulses applied to all qubits of the same color $c$, i.e., by $\widetilde{X}_{c}\equiv\bigotimes_{v\in V_{c}}X_{v}$, we flip the sign of all $\sigma^{\alpha}_{v}$ terms corresponding to qubits $v$ of color $c$ since they anticommute with $\widetilde{X}_{c}$. Similarly, for the two body terms, we flip the sign of all $\sigma^{\alpha}_{u}\sigma^{\beta}_{v}$ terms where one of the qubits $u,v$ is of color $c$ and the corresponding Pauli operator anticommutes with $X$. Thus,
$\widetilde{X}_{c}f_{\tau}\widetilde{X}_{c}^{\dagger}=\exp\left[-i\tau\left(\widetilde{H}_{1,c}+\widetilde{H}_{2,c}\right)\right]$ (3a)
$\widetilde{H}_{1,c}=\sum_{\alpha\in\{x,y,z\}}\widetilde{H}^{\alpha}_{1,c}\ ,\quad\widetilde{H}_{2,c}=\sum_{\alpha,\beta\in\{x,y,z\}}\widetilde{H}^{\alpha\beta}_{2,c}$ (3b)
$\widetilde{H}^{\alpha}_{1,c}=\widetilde{X}_{c}H^{\alpha}_{1}\widetilde{X}_{c}^{\dagger}=\sum_{c^{\prime}\in C}(-1)^{\delta_{cc^{\prime}}(1-\delta_{\alpha x})}\sum_{v\in V_{c^{\prime}}}\sigma^{\alpha}_{v}$ (3c)
$\widetilde{H}^{\alpha\beta}_{2,c}=\widetilde{X}_{c}H^{\alpha\beta}_{2}\widetilde{X}_{c}^{\dagger}=$ (3d)
$\quad\sum_{c_{1}\neq c_{2}}(-1)^{\delta_{cc_{1}}(1-\delta_{\alpha x})}(-1)^{\delta_{cc_{2}}(1-\delta_{\beta x})}\!\!\!\!\sum_{(u,v)\in E_{c_{1},c_{2}}}\!\!\!\!\sigma^{\alpha}_{u}\sigma^{\beta}_{v}.$
Note that $\widetilde{H}^{x}_{1,c}=H^{x}_{1}$ and $\widetilde{H}^{xx}_{2,c}=H^{xx}_{2}$. To go beyond qubits of a single color $c$, we form a criterion for determining whether to conjugate $f_{\tau}$ by $\widetilde{X}_{c}$ for a given color $c$ at time step $j$ according to the Hadamard matrix.
Let $\nu=\left\lfloor\log_{2}\chi\right\rfloor+1$, $N=2^{\nu}$, and consider the $N\times N$ Hadamard matrix
$W_{\nu}\equiv W_{1}^{\otimes\nu}=\frac{1}{\sqrt{N}}\sum_{i,j\in\{0,1\}^{\nu}}(-1)^{i\cdot j}|{i}\rangle\!\langle j|,$ (4)
where $W_{1}$ is the standard $2\times 2$ Hadamard matrix and where the dot product is the bit-wise scalar product with addition modulo 2. Below, we use $i$ and $j$ to denote an integer or its binary expansion (i.e., $i=i_{0}i_{i}\dots i_{\nu-1}$), depending on the context.
Consider any injective function $g:C\rightarrow\{1,2,\dots,N-1\}$ that maps distinct colors in $C$ to distinct rows $i>0$ of the Hadamard matrix $W_{\nu}$. Note that $\nu$ is the number of bits needed to account for every color $c\in C$. For every such color, we use row $i=g(c)$ to schedule the decoupling scheme for the qubits in $V_{c}$: at each time step $j=0,1,\dots,N-1$, if $\bra{i}W_{\nu}\ket{j}=(-1)^{i\cdot j}=-1$, i.e., if $i\cdot j\equiv 1\mod 2$, we conjugate $f_{\tau}$ by $\widetilde{X}_{c}$. Equivalently, for each color $c$ at each time step $j$, we conjugate $f_{\tau}$ by $\widetilde{X}_{c}^{g(c)\cdot j}$. The resulting unitary control propagator across all colors $c\in C$ at time step $j$ is then
$U^{x}_{j}\equiv\overline{X}_{j}f_{\tau}\overline{X}_{j}^{\dagger}\ ,\quad\overline{X}_{j}\equiv\bigotimes_{c\in C}\widetilde{X}_{c}^{g(c)\cdot j}\ .$ (5)
The only difference from Eq. 3 is that now the indicator functions are generalized from a single color (such as $\delta_{cc^{\prime}}$) to the set of all colors dictated by the Hadamard matrix at time step $j$, i.e., to $g(c)\cdot j$. Consequently, Eq. 3 is replaced by
$U^{x}_{j}=\exp\left[-i\tau\left(\overline{H}_{1,j}+\overline{H}_{2,j}\right)\right]$ (6a)
$\overline{H}_{1,j}=\sum_{\alpha\in\{x,y,z\}}\overline{H}^{\alpha}_{1,j}\ ,\ \ \ \overline{H}_{2,j}=\sum_{\alpha,\beta\in\{x,y,z\}}\overline{H}^{\alpha\beta}_{2,j}$ (6b)
$\overline{H}^{\alpha}_{1,j}=\overline{X}_{j}H^{\alpha}_{1}\overline{X}_{j}^{\dagger}=\sum_{c\in C}(-1)^{g(c)\cdot j(1-\delta_{\alpha x})}\sum_{v\in V_{c}}\sigma^{\alpha}_{v}$ (6c)
$\overline{H}^{\alpha\beta}_{2,j}=\overline{X}_{j}H^{\alpha\beta}_{2}\overline{X}_{j}^{\dagger}=$ (6d)
$\quad\sum_{c_{1}\neq c_{2}}(-1)^{g(c_{1})\cdot j(1-\delta_{\alpha x})}(-1)^{g(c_{2})\cdot j(1-\delta_{\beta x})}\!\!\!\!\sum_{(u,v)\in E_{c_{1},c_{2}}}\!\!\!\!\sigma^{\alpha}_{u}\sigma^{\beta}_{v}.$
Similarly, $\overline{H}^{x}_{1,j}=H^{x}_{1}$ and $\overline{H}^{xx}_{2,j}=H^{xx}_{2}$, meaning that pure-$x$ terms are invariant.
Each $U^{x}_{j}$ takes one time step of length $\tau$. Then, using the BCH expansion, the entire sequence of $N$ time steps and total duration $T=N\tau$ has the following unitary:
$\prod_{j=0}^{N-1}U^{x}_{j}$ $=\exp\Big{[}-\!i\tau\sum_{j=0}^{N-1}\overline{H}_{1,j}+\overline{H}_{2,j}\Big{]}+\mathcal{O}\left(T^{2}\right).$ (7)
Now, observe that the single-qubit sum $\sum_{j}\overline{H}_{1,j}=NH^{x}_{1}$, since $\sum_{j=0}^{N-1}(-1)^{g(c)\cdot j}=0$ $\forall~{}g(c)\neq 0$, where we used the fact that all rows $i=g(c)>0$ of the Hadamard matrix have entries that add up to zero. Similarly, the two-qubit sum $\sum_{j}\overline{H}_{2,j}=NH^{xx}_{2}$, since $\sum_{j=0}^{N-1}(-1)^{g(c_{1})\cdot j}(-1)^{g(c_{2})\cdot j}=\delta_{g(c_{1}),g(c_{2})}=0$, where we used the fact that the Hadamard matrix $W_{\nu}$ is orthogonal, i.e., $\delta_{i_{1},i_{2}}=\bra{i_{1}}W_{\nu}^{T}W_{\nu}\ket{i_{2}}=\frac{1}{N}\sum_{j\in\{0,1\}^{\nu}}(-1)^{i_{1}\cdot j}(-1)^{i_{2}\cdot j}$, and that $g$ is injective, i.e., $c_{1}\neq c_{2}\implies i_{1}=g(c_{1})\neq g(c_{2})=i_{2}$. Thus,
$U^{x}\equiv\prod_{j=0}^{N-1}U^{x}_{j}=\exp\left[-iT(H^{x}_{1}+H^{xx}_{2})\right]+\mathcal{O}\left(T^{2}\right),$ (8)
and we have eliminated to first order in $\tau$ all non-pure-$x$ terms in $H_{1}$ and $H_{2}$. The total number of time steps, i.e., circuit depth, is $N=2^{\nu}=2^{\left\lfloor\log_{2}\chi\right\rfloor+1}$, i.e., $\chi\leq N\leq 2\chi$. ∎
Definition 1.
The sequence $U^{x}=\prod_{j=0}^{N-1}U^{x}_{j}$ is called single-axis $x$-type CHaDD. Other single-axis CHaDD sequences are obtained by replacing $x$ with another axis.
As is the case in the single-qubit XY4 sequence [11], we can obtain multi-axis CHaDD sequences by concatenating single-axis CHaDD sequences about perpendicular axes [35]. This leads a concatenated multi-axis CHaDD protocol requiring a quadratic circuit depth in $\chi(G)$; see the Appendix. However, as we show next, for $\chi(G)\geq 3$ a more efficient multi-axis CHaDD protocol is possible, that retains the linear scaling in $\chi(G)$. For $\chi(G)=2$, both protocols require a circuit depth of $16$.
[…]
Theorem 2 (multi-axis CHaDD).
Under the same assumptions as in 1, a circuit depth of $3\chi+1\leq N\leq 2(3\chi+5)$, involving only single-qubit Pauli pulses, suffices to cancel to first order in $\tau$ all terms in $H$ on a qubit connectivity graph $G$ with chromatic number $\chi$.
The proof combines ideas from Refs. [57, 61] with our single-axis CHaDD approach.
Proof.
Assign to each color $c$ a Schur subset via an injective function $h$. For a given Schur subset $h(c)$, label its three rows $h(c;x)$, $h(c;y)$, and $h(c;z)$. Define the $\chi\times N$ sign matrix $S_{\alpha}$, where $\alpha\in\{x,y,z\}$, as the matrix whose rows are $\{h(c;\alpha)\}_{c\in C}$. These sign matrices $S_{\alpha}$ are Schur-closed. This implies that for each fixed $(c,j)$’th entry, $([S_{x}]_{cj},[S_{y}]_{cj},[S_{z}]_{cj})$ can only be one of (+,+,+)(+,+,+)( + , + , + ), (+,,)(+,-,-)( + , - , - ), (,+,)(-,+,-)( - , + , - ), or (,,+)(-,-,+)( - , - , + ). These tuples correspond to whether or not a given Pauli operator $\widetilde{\sigma}_{(c,j)}$ acting on all qubits of color $c$ commutes (+) or anticommutes (-) with $(\widetilde{X}_{c},\widetilde{Y}_{c},\widetilde{Z}_{c})$, i.e., $\widetilde{\sigma}_{(c,j)}=\widetilde{I}_{c}$, $\widetilde{X}_{c}$, $\widetilde{Y}_{c}$, and $\widetilde{Z}_{c}$, respectively. The signs acquired after conjugating Pauli terms $(\sigma^{\alpha}_{v})_{\alpha\in\{x,y,z\},f(v)=c}$ by $\widetilde{\sigma}_{(c,j)}$ are the $(c,j)$’th entries $([S_{\alpha}]_{cj})_{\alpha\in\{x,y,z\}}$.
Then the unitary control propagator is $\overline{\sigma}_{j}\equiv\bigotimes_{c\in C}\widetilde{\sigma}_{(c,j)}$, and when we conjugate $f_{\tau}$ by $\overline{\sigma}_{j}$, i.e., $U_{j}\equiv\overline{\sigma}_{j}f_{\tau}\overline{\sigma}_{j}^{\dagger}$, the sign acquired by $\sigma_{v}^{\alpha}$ is $[S_{\alpha}]_{cj}$ if $f(v)=c$, while the sign acquired by $\sigma_{u}^{\alpha}\sigma_{v}^{\beta}$ ($u<v$) is $[S_{\alpha}]_{c_{1}j}[S_{\beta}]_{c_{2}j}$ if $f(u)=c_{1}$ and $f(v)=c_{2}$:
$U_{j}=\exp\left[-i\tau\left(\overline{H}_{1,j}+\overline{H}_{2,j}\right)\right]$ (9a)
$\overline{H}_{1,j}=\sum_{\alpha\in\{x,y,z\}}\overline{H}^{\alpha}_{1,j}\ ,\ \ \ \overline{H}_{2,j}=\sum_{\alpha,\beta\in\{x,y,z\}}\overline{H}^{\alpha\beta}_{2,j}$ (9b)
$\overline{H}^{\alpha}_{1,j}=\overline{\sigma}_{j}H^{\alpha}_{1}\overline{\sigma}_{j}^{\dagger}=\sum_{c\in C}[S_{\alpha}]_{cj}\sum_{v\in V_{c}}\sigma^{\alpha}_{v}$ (9c)
$\overline{H}^{\alpha\beta}_{2,j}=\overline{\sigma}_{j}H^{\alpha\beta}_{2}\overline{\sigma}_{j}^{\dagger}=$ (9d)
$\quad\sum_{c_{1}\neq c_{2}}[S_{\alpha}]_{c_{1}j}[S_{\beta}]_{c_{2}j}\!\!\!\!\sum_{(u,v)\in E_{c_{1},c_{2}}}\!\!\!\!\sigma^{\alpha}_{u}\sigma^{\beta}_{v}.$
Then the overall unitary for all $N$ time steps is
$U\equiv\prod_{j=0}^{N-1}U_{j}$ $=\exp\Big{[}-\!i\tau\sum_{j=0}^{N-1}\overline{H}_{1,j}+\overline{H}_{2,j}\Big{]}+\mathcal{O}\left(T^{2}\right).$ (10)
To remove all one- and two-local terms, i.e., $\sum_{j}\overline{H}_{1,j}=0$ and $\sum_{j}\overline{H}_{2,j}=0$, we need $\sum_{j}[S_{\alpha}]_{cj}=0$ and $\sum_{j}[S_{\alpha}]_{c_{1}j}[S_{\beta}]_{c_{2}j}=0$. The first condition follows immediately from the fact that these are rows of the Hadamard matrix, and the second follows from the fact that each Schur-subset was drawn from different partitions ($c_{1}\neq c_{2}$) of the Hadamard matrix, so the rows of the sign matrices $S_{\alpha}$ are orthogonal.
The Hadamard matrix $W_{\nu}$ can be partitioned into exactly $S_{e}=(2^{\nu}-1)/3$ Schur-subsets if $\nu$ is even and at most $S_{o}=(2^{\nu}-5)/3$ if $\nu$ is odd [57] (see also [73, Thm. 4.1] and the Appendix). Let $\nu=2k$ and $k\in\mathbb{Z}^{+}$, so $S_{e}(k)=(2^{2k}-1)/3$ and $S_{o}=(2^{2k+1}-5)/3$. Since the circuit depth is $N=2^{\nu}$, we would like to minimize $k$. We need at least as many Schur subsets as colors, hence $\min\{S_{e}(k),S_{o}(k)\}\geq\chi$. Minimizing $k$, we have $k=\min\{f_{e}(\chi),f_{o}(\chi)\}$, where $f_{e}(\chi)=\lceil\frac{1}{2}\log_{2}(3\chi+1)\rceil$ and $f_{o}(\chi)=\lceil\frac{1}{2}[\log_{2}(3\chi+5)-1]\rceil$. Thus, we obtain $N=\min\{2^{2\lceil\frac{1}{2}\log_{2}(3\chi+1)\rceil},2^{2\lceil\frac{1}{2}[\log_{2}(3\chi+5)-1]\rceil+1}\}\in\Theta(\chi)$, i.e., $3\chi+1\leq N\leq 2(3\chi+5)$. ∎
We explain the upper and lower bounds in the Appendix. Since $N$ is a step function in $\chi$, these bounds are loose (e.g., $N=16$ for $2\leq\chi\leq 5$, and $N=32$ for $6\leq\chi\leq 9$) but they are nearly tight envelopes.
Examples.—We illustrate single-axis $x$-type CHaDD for a few examples with low chromatic numbers. Since $\chi(G)\geq 2$ for graphs with $|E|\geq 1$, we use graphs $G$ with $\chi(G)=2,3$, i.e., $\nu=\left\lfloor\log_{2}\chi(G)\right\rfloor+1=2$. By 1, a circuit depth of $N=2^{\nu}=4$ is then sufficient to cancel all non-pure-$x$ interactions (e.g., $ZZ$ crosstalk) and qubit decoherence to first order, scheduled according to the Hadamard matrix $W_{2}$ shown in Fig. 1 (middle).
Figure 1: Color-to-row maps for $2$-colorable (left) and $3$-colorable (right) graphs. Solid arrows correspond to an unbalanced schedule, dashed to a balanced schedule. Middle: the Hadamard matrix $W_{2}$.
Per 1, the first row ($i=0$) is not used. For $\chi(G)=3$, we must use all three of the remaining rows $i=1,2,3$, but for $\chi(G)=2$, we can choose any pair of rows.
Example 1: $W_{2}$ for $\chi(G)=2$.—Consider a square grid of qubits with nearest-neighbor coupling, as in Fig. 2 (top and middle). This graph $G$ is $2$-colorable (bipartite), i.e., there exists a proper $2$-coloring $f:V\rightarrow\{1,2\}\equiv C$. Examples of such graphs include the Rigetti Ankaa family of chips [74] and the IBM heavy-hex layout [75].
Let us choose the identity color-to-row map $i=g(c)=c$ so that colors $c=1,2$ correspond to rows $i=1,2$, respectively, as in Fig. 1 (left, solid arrows). We conjugate the free evolution $f_{\tau}$ by all $\widetilde{X}_{i}$ operators for which $i\right|W_{2}\left|j\right>=-1$ for each time step (column) $j$. This readily yields $U^{x}_{0}=f_{\tau}$ for the first interval, $U^{x}_{1}=\widetilde{X}_{1}f_{\tau}\widetilde{X}_{1}^{\dagger}$ for the second, $U^{x}_{2}=\widetilde{X}_{2}f_{\tau}\widetilde{X}_{2}^{\dagger}$ for the third, and, since $\overline{X}_{3}=\widetilde{X}_{1}\widetilde{X}_{2}$, finally $U^{x}_{3}=(\widetilde{X}_{1}\widetilde{X}_{2})f_{\tau}(\widetilde{X}_{1}\widetilde{X}_{2})^{\dagger}$ for the last interval. Taking the product of all unitaries yields the unbalanced schedule
$U^{x}=\prod_{j=0}^{3}U^{x}_{j}$ $=\widetilde{X}_{1}\widetilde{X}_{2}f_{\tau}\widetilde{X}_{1}f_{\tau}\widetilde{X}_{1}\widetilde{X}_{2}f_{\tau}\widetilde{X}_{1}f_{\tau},$ (11)
wherein a pulse is applied to all $1$-colored qubits every $\tau$ but only every $2\tau$ to the $2$-colored qubits; see Fig. 2 (top).
A balanced schedule results if instead we choose $i=g(c)=c+1$, as in Fig. 1 (left, dashed arrows), resulting in $U^{x}_{0}=f_{\tau}$, $U^{x}_{1}=\widetilde{X}_{2}f_{\tau}\widetilde{X}_{2}^{\dagger}$, $U^{x}_{2}=(\widetilde{X}_{1}\widetilde{X}_{2})f_{\tau}(\widetilde{X}_{1}\widetilde{X}_{2})^{\dagger}$, and $U^{x}_{3}=\widetilde{X}_{1}f_{\tau}\widetilde{X}_{1}^{\dagger}$, so that
$U^{x}=\prod_{j=0}^{3}U^{x}_{j}=\widetilde{X}_{1}f_{\tau}\widetilde{X}_{2}f_{\tau}\widetilde{X}_{1}f_{\tau}\widetilde{X}_{2}f_{\tau},$ (12)
wherein a pulse is applied alternately to $1$ and $2$-colored qubits every $2\tau$. This schedule is illustrated in Fig. 2 (middle) and is preferred due to its symmetry and lower overall pulse count of four single-color pulses vs six for the unbalanced schedule. It is an interesting question to identify the general conditions for a balanced schedule directly from the Hadamard matrix in arbitrary dimensions.
Figure 2: Examples of schedules for 2-colorable (top and middle) and 3-colorable (bottom) graphs. Purple circles indicate DD pulses, and time flows from left to right. Top row: unbalanced schedule for a $2$-colorable graph [Eq. 11]. Middle row: balanced schedule for a $2$-colorable graph [Eq. 12]. Bottom row: unbalanced schedule for a $3$-colorable graph [Eq. 13].
Example 2: $W_{2}$ for $\chi(G)=3$.—Examples for which $\chi(G)=3$ are hexagonal or triangular grids [Fig. 2 (bottom)], whose graph degree is also $3$. We must now use all three rows $i=1,2,3$ of the Hadamard matrix $W_{2}$, and $g$ can only be a permutation. Consider the identity: $i=g(c)=c$, as in Fig. 1 (right). We obtain $U^{x}_{0}=f_{\tau}$, $\overline{X}_{1}=\widetilde{X}_{1}\widetilde{X}_{3}$ so that $U^{x}_{1}=(\widetilde{X}_{1}\widetilde{X}_{3})f_{\tau}(\widetilde{X}_{1}\widetilde{X}_{3})^{\dagger}$, and similarly $U^{x}_{2}=(\widetilde{X}_{2}\widetilde{X}_{3})f_{\tau}(\widetilde{X}_{2}\widetilde{X}_{3})^{\dagger}$ and $U^{x}_{3}=(\widetilde{X}_{1}\widetilde{X}_{2})f_{\tau}(\widetilde{X}_{1}\widetilde{X}_{2})^{\dagger}$. Thus,
$U^{x}=\prod_{j=0}^{3}U^{x}_{j}=\widetilde{X}_{1}\widetilde{X}_{2}f_{\tau}\widetilde{X}_{1}\widetilde{X}_{3}f_{\tau}\widetilde{X}_{1}\widetilde{X}_{2}f_{\tau}\widetilde{X}_{1}\widetilde{X}_{3}f_{\tau},$ (13)
an unbalanced schedule using eight single-color pulses; see Fig. 2 (bottom). A balanced schedule is impossible in this case using only six single-color pulses.
Conclusions and outlook.—We have proposed CHaDD: a first-order DD scheme that is efficient in the chromatic number of the qubit connectivity graph, which leads to a significant reduction in the circuit depth required to decouple arbitrary connectivity graphs relative to previously known multi-qubit decoupling methods [57, 58, 59, 60, 61, 62, 63, 20, 21, 22]. This includes, as a special case, multi-qubit crosstalk decoupling.
To generalize CHaDD beyond first-order decoupling is straightforward. For example, concatenation of multi-axis CHaDD with itself is the direct multi-qubit generalization of concatenating XY4 with itself, which leads to the single-qubit concatenated DD (CDD) family [35], yielding an extra suppression order in time-dependent perturbation theory with every additional level of concatenation [36, 50].
However, concatenation incurs an exponential cost in circuit depth, and it would be desirable to be able to replace the uniform pulse intervals used in the construction of single-axis CHaDD with the non-uniform intervals used in the single-qubit Uhrig DD (UDD) sequence family [37], in order to guarantee the same higher-order performance as UDD [39] in the single-axis, multi-qubit setting. This would directly extend to multi-axis single-qubit decoupling via the quadratic DD (QDD) sequence [42], which requires at most $(m+1)^{2}$ pulses for order-$m$ suppression in terms of the Dyson series expansion [43]. Successfully combining CHaDD with QDD would generally be much more efficient than the multi-qubit, nested UDD (NUDD) sequence, which scales exponentially with the number of qubits [41]. Whether these extensions of CHaDD are possible is an interesting open problem.
It should be possible to generalize CHaDD beyond qubits to multi-level systems by replacing the Pauli matrices used in the proof of 1 with elements of the generalized Pauli (or Heisenberg-Weyl) group [76, 77, 60]. Finally, it is an interesting open problem to generalize CHaDD from two-local and instantaneous pulses to $k$-local interactions and bounded controls for $n$ qubits, for which the current state of the art is $\mathcal{O}[n^{k-1}\log(n)]$ [62].
Acknowledgments.—This research was supported by the Army Research Office MURI grant W911NF-22-S-0007 and by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA) and the Army Research Office, under the Entangled Logical Qubits program through Cooperative Agreement Number W911NF-23-2-0216. The authors gratefully acknowledge Victor Kasatkin and Jenia Mozgunov for helpful discussions.
Appendix A Multi-axis CHaDD via concatenation
Corollary A1 (Concatenated multi-axis CHaDD).
Assume the free evolution unitary is $f_{\tau}=e^{-i\tau H}$, where $H$ is given by Eq. 1. Then a circuit depth of $4^{\left\lfloor\log_{2}\chi(G)\right\rfloor+1}\leq 4\chi^{2}(G)$ is sufficient to completely cancel $H$ to first order in $\tau$ on a qubit connectivity graph $G$ with chromatic number $\chi(G)$.
Proof.
It follows immediately from 1 (by swapping the $x$ and $z$ indices) that a single-axis $z$-type CHaDD sequence has an overall unitary of
$U^{z}\equiv\prod_{j=0}^{N-1}U^{z}_{j}=\exp\left[-iT(H^{z}_{1}+H^{zz}_{2})\right]+\mathcal{O}\left(T^{2}\right).$ (14)
If we concatenate the single-axis $x$- and $z$-type sequences at every time step, i.e., replace $f_{\tau}$ in $U^{x}_{j}=\overline{X}_{j}f_{\tau}\overline{X}_{j}^{\dagger}$ with the single-axis $z$-type CHaDD sequence $U^{z}$, we obtain $U^{xz}_{j}\equiv\overline{X}_{j}U^{z}\overline{X}_{j}^{\dagger}$, so that
$U^{xz}\equiv\prod_{j=0}^{N-1}U^{xz}_{j}=\prod_{j=0}^{N-1}\overline{X}_{j}e^{-iT(H^{z}_{1}+H^{zz}_{2})}\overline{X}_{j}^{\dagger}+\mathcal{O}\left(NT^{2}\right).$ (15)
This is just the single-axis $x$-type CHaDD sequence $U^{x}$ given by Eq. 8 applied to an effective Hamiltonian $H=H^{z}_{1}+H^{zz}_{2}$ without any $H_{1}^{x}$ or $H_{2}^{xx}$ terms, so it follows from 1 that this Hamiltonian is completely canceled to first order, i.e.,
$\prod_{j=0}^{N-1}\overline{X}_{j}e^{-iT(H^{z}_{1}+H^{zz}_{2})}\overline{X}_{j}^{\dagger}=I+\mathcal{O}\left(NT^{2}\right).$ (16)
Hence, the total CHaDD unitary is $U^{xz}=I+\mathcal{O}\left(NT^{2}\right)$. This requires a circuit depth of $N^{2}=2^{2(\left\lfloor\log_{2}\chi(G)\right\rfloor+1)}\leq 4\chi^{2}(G)$. ∎
Definition A1.
The sequence $U^{xz}=\prod_{j=0}^{N-1}U^{xz}_{j}$ is called concatenated multi-axis CHaDD.
Clearly, replacing $x$ and $z$ with other combinations of orthogonal axes is equivalent to $U^{xz}$. Thus, $U^{xz}$ can be interpreted as the efficient multi-axis, multi-qubit generalization of the XY4 sequence [11].
Appendix B Connections to projective geometry
The properties of Schur subsets of a sign matrix presented by Leung [57] were discovered in a more general form several decades earlier in the projective geometry community, using the terminology of partial $t$-spreads in the Desargesian projective space $PG(d,q)$, where $d$ is the dimension and $q$ is the order (number of elements) of the finite field $\mathbb{F}_{q}$ [73]. In such a space, points and lines are defined in terms of vector spaces over $\mathbb{F}_{q}$, and the incidence structure obeys the axioms of projective geometry. Desargues’ theorem guarantees that any configuration of points and lines that forms a perspective triangle has collinear points, ensuring a well-defined projective structure. A perspective triangle is a configuration where two triangles are in perspective from a point or a line
For our purposes, $PG(d,q)=\mathbb{Z}_{2}^{\nu}\setminus\{0\}$, $d=\nu-1$, $q=2$, $t=1$, and the size of the partial $t$-spread is the number of Schur subsets.
In this context, Ref. [73, Thm. 4.1] states that there are no more than $(2^{\nu}-5)/3$ Schur subsets for a Hadamard matrix $W_{\nu}$ with odd $\nu$. For even $\nu$ one can pick all $12^{\nu}-$ 1 non-trivial rows of the Hadamard matrix (one row is the trivial row of all $+$), resulting in $(2^{\nu}-1)/3$ Schur subsets, so the upper bound is not needed.
Moreover, the bounds for both odd and even $\nu$ are achievable: for even $\nu$ Ref. [73] states that this was known before (Results 2.1 and 2.2), and for odd $\nu$ this is shown in Theorem 4.2.
Appendix C Circuit depth of multi-axis CHaDD
In the main text, we showed that
$N=\min\{2^{2\lceil\frac{1}{2}\log_{2}(3\chi+1)\rceil},2^{2\lceil\frac{1}{2}[\log_{2}(3\chi+5)-1]\rceil+1}\}$ (17)
for the circuit depth of multi-axis CHaDD. Here we derive the upper and lower bounds. It follows from the definitions of $S_{e}$ and $S_{o}$ that $12^{\nu}\geq 3\chi+$ 1 and $52^{\nu}\geq 3\chi+$ 5 in the case that $\nu$ is even and odd, respectively. Thus, in either case, we have that $N=2^{\nu}\geq 3\chi+1$, establishing a lower bound on the circuit depth regardless of the value of $\nu$. Now consider that we wish to choose the smallest value of $\nu$ that satisfies the above inequalities so that we may state $2^{\nu}\geq 3\chi+1\geq 2^{\nu-1}$ and $2^{\nu}\geq 3\chi+5\geq 2^{\nu-1}$ if $\nu$ is even or odd, respectively, from which it follows that $2(3\chi+5)\geq 2^{\nu}=N$. Hence, $12(3\chi+5)\geq N\geq 3\chi+$ 1. These results are shown in Fig. 3.
Figure 3: Circuit depth $N(\chi)$ of multi-axis CHaDD [Eq. 17] along with the upper and lower bounds, $2(3\chi+5)$ and $13\chi+$ 1, respectively.